# ELUTQ: Efficient LUT-Aware Quantization for Deploying Large Language Models on Edge Devices

### v1

@@ Abstract
The deployment of Large Language Models on CPU-based edge devices is crucial for enabling on-device intelligence and broadening the reach of AI applications. However, such deployment remains challenging due to the limited memory and computational resources typical of these devices. When performing inference on edge CPUs, memory usage and latency are two primary bottlenecks. Weight quantization effectively reduces memory consumption, yet existing hardware-friendly methods often rely on uniform quantization, which suffers from poor weight-distribution fitting and high dequantization overhead under low-bit settings. To address these issues, we propose ELUTQ, an efficient quantization framework featuring a novel quantization format termed Hierarchical Linear Quantization (HLQ). HLQ is designed to better capture the statistical characteristics of weights without increasing the computational cost of Bit-serial LUT-based GEMM operations, thereby eliminating dequantization overhead. As a fundamental quantization scheme, HLQ is orthogonal to existing quantization algorithms and can be seamlessly integrated into various quantization pipelines. To enable efficient deployment on edge devices, ELUTQ designs high-performance CPU kernels to support end-to-end inference. Extensive experiments demonstrate the effectiveness of our approach. For the LLaMA3-8B model, when combined with the post-training quantization framework, HLQ enhances uniform quantization by achieving approximately 8% perplexity reduction at 3-bit precision and 85% perplexity reduction at 2-bit precision, with quantization completed within one hour. When combined with efficient finetuning techniques, HLQ further improves perplexity under the 2-bit configuration and completes quantization in about two hours. In terms of inference efficiency, under a 4-thread configuration with batch size = 1, our 2-bit quantized LLaMA2-7B model achieves a throughput of over 25 tokens per second on an Apple M2 chip. All the code is available at https://github.com/Nkniexin/ELUTQ.
@@ 1 Introduction
Many large model quantization methods aim to efficiently complete model quantization, where efficiency primarily refers to time and memory consumption, particularly video memory (VRAM), which is typically more constrained and expensive than CPU memory. Achieving quantization with low memory usage and short processing time is critical, as it enables running quantization algorithms even on consumer-grade GPUs such as the RTX 3090 or RTX 4090. Most post-training quantization approaches (Frantar et al., 2022; Lin et al., 2024; Xiao et al., 2023; Huang et al., 2024; Shang et al., 2023; Zhao & Yuan, 2025; Kim et al., 2023) can complete quantization quickly and with modest memory requirements. These train-free methods operate without retraining and typically quantize the model layer by layer. For example, GPTQ quantizes one linear layer at a time. With sufficient CPU memory, even large models such as LLaMA2-70B can be quantized on a single A6000 GPU. In contrast, efficient finetuning–based methods (Shao et al., 2023; Dettmers et al., 2023b, a; Chee et al., 2023; Li et al., 2023; Xu et al., 2023; Chen et al., 2024b) require partial retraining, which increases time cost but often mitigates VRAM pressure by freezing most parameters, training only a small subset, or performing quantization in batches. This strategy allows quantization to be completed within limited video memory. In this paper, we combine our proposed Hierarchical Linear Quantization with efficient quantization techniques to enhance model accuracy while maintaining low VRAM consumption and short quantization time, thereby further advancing on-device intelligence.
Although weight quantization reduces memory footprint, traditional methods require dequantization to higher precision (8-bit/FP16) for computation, introducing significant overhead that can paradoxically slow down inference at low bit-widths. Recent advances (Wei et al., 2025; Park et al., 2025, 2022) address this by replacing standard GEMM with lookup table (LUT)-based operations that implement generalized FP-INT multiplication, eliminating dequantization while achieving both linear latency reduction with bit-width and improved energy efficiency. FLGLUT (Park et al., 2025) accelerates these operations on GPUs and T-MAC (Wei et al., 2025) optimizes for CPUs via SIMD instructions. We enhance this paradigm through a pure C++ kernel redesign that specifically supports our novel Hierarchical Linear Quantization format, maintaining cross-platform compatibility while optimizing for our method’s unique requirements.
@@ 2.2 LUT-Based GEMM
Figure 2(c) illustrates another GEMM paradigm termed Bit-serial LUT-based GEMM. Unlike the scheme in Figure 2(b), this method utilizes the LUT to store all possible dot products between an activation vector and single-bit weights. The detail computational procedure is shown in Figure 3. First, a $q$-bit quantized weight matrix $\mathrm{W}_{int}$ is decomposed into $q$ single-bit matrices $\{W_{0},W_{1},...,W_{q-1}\}$ offline, where each element is either 0 or 1, representing the respective bit planes of the original weights. For example, for integer values (9, 7, 6, 3) with binary representations (1001, 0111, 0110, 0011), the matrix for the lowest bit is (1, 1, 0, 1), and the matrix for the highest bit is (1, 0, 0, 0). This decomposition is performed offline, incurring no runtime overhead. During inference, for an activation vector of the group size $g$, the system precomputes the dot products between this activation vector and all $2^{g}$ possible combinations of single-bit weights, storing the results in the LUT. Thus, the original matrix computation requiring high-precision multiply-accumulate operations is simplified into highly efficient table lookups followed by summation. This paradigm has been demonstrated to offer high computational efficiency and energy efficiency (Park et al., 2025; Wei et al., 2025). For example, FIGLUT (Park et al., 2025) optimized the table structure for GPU architectures to avoid bank conflicts, while T-MAC leveraged CPU vectorized lookup instructions (AVX2/NEON) to enable efficient LUT operations on CPUs.
@@ 3.1 Bell-Shaped Distribution of Weights
As noted in prior studies (Kim et al., 2023; Dettmers et al., 2023a; Zhao & Yuan, 2025), weights often exhibit a bell-shaped distribution. Uniform quantization, which maps weights to a uniformly spaced grid, struggles to accurately capture such distributions, especially under low-bit settings. Figure 4 shows that the weight distribution of an out_channel in the up_projection layer of the LLaMA3-8B model closely follows a Gaussian pattern, with values heavily concentrated near zero. This characteristic, commonly observed across various models, highlights the inherent limitation of uniform quantization in effectively representing natural weight distributions. Consequently, non-uniform quantization strategies are often necessary to achieve higher fidelity in low-bit compression.
Figure 4: (Top) The weight distribution of one output channel in a up_proj of LLaMA3-8B. (Bottom) MSE loss for weight comparsion between Uniform and HLQ.
@@ 3.2 Optimization for Bit-Serial LUT-Based GEMM
For ordinary LUT-based GEMM algorithms shown in figure 2(b), previous efforts such as SqueezeLLM (Kim et al., 2023) and GANQ (Zhao & Yuan, 2025) have achieved significant improvements in algorithmic performance for 3-bit and 4-bit quantization, along with the development of highly efficient CUDA kernels for accelerated inference. However, these implementations still lack support for CPU devices. Regarding Bit-serial LUT-Based GEMM, existing works including T-MAC (Wei et al., 2025), LUT-GEMM (Park et al., 2022), and FIGULT (Park et al., 2025) have primarily focused on optimizing kernel design to support efficient uniform quantization schemes, yet they do not address algorithmic enhancements for improving quantization accuracy. This limitation restricts the broader application of such kernels. To bridge this gap, our work proposes using HLQ to boost quantization accuracy at the algorithm level, while leveraging Bit-Serial LUT-Based GEMM to enable efficient inference on CPU devices.

### v2

@@ Abstract
Weight quantization effectively reduces memory consumption and enable the deployment of Large Language Models on CPU-based edge devices, yet existing hardware-friendly methods often rely on uniform quantization, which suffers from poor weight-distribution fitting and high dequantization overhead under low-bit settings. In this paper, we propose ELUTQ, an efficient quantization framework featuring a novel quantization format termed Hierarchical Linear Quantization (HLQ). HLQ is designed to better capture the statistical characteristics of weights without increasing the computational cost of Bit-serial LUT-based GEMM operations, thereby eliminating dequantization overhead. HLQ is orthogonal to existing quantization algorithms. For the LLaMA3.1-8B model, when combined with the post-training quantization framework, HLQ enhances uniform quantization by achieving approximately 8% perplexity reduction at 3-bit precision and 85% perplexity reduction at 2-bit precision. When combined with efficient finetuning techniques, HLQ further improves model accuracy. A novel disk-offload technique is integrated into ELUTQ, enabling it to complete the quantization of LLaMA 3.1–70B using only 64 GB of CPU memory and 48 GB of VRAM, significantly reducing the hardware requirements for large-scale model quantization. To enable efficient deployment on edge devices, ELUTQ designs high-performance CPU kernels to support end-to-end inference. Under a 4-thread configuration with batch size = 1, our 2-bit quantized LLaMA2-7B model achieves a throughput of over 25 tokens per second on an Apple M2 chip. All the code is available at https://github.com/Nkniexin/ELUTQ.
@@ 1 Introduction
Many large model quantization methods aim to efficiently complete model quantization, where efficiency primarily refers to time and memory consumption, particularly video memory (VRAM), which is typically more constrained and expensive than CPU memory. Achieving quantization with low memory usage and short processing time is critical, as it enables running quantization algorithms even on consumer-grade GPUs such as the RTX 3090 or RTX 4090. Most post-training quantization approaches (Frantar et al., 2022; Lin et al., 2024; Xiao et al., 2023; Huang et al., 2024; Shang et al., 2023; Zhao & Yuan, 2025; Kim et al., 2023) can complete quantization quickly and with modest memory requirements. These train-free methods operate without retraining and typically quantize the model layer by layer. For example, GPTQ quantizes one linear layer at a time. With sufficient CPU memory, even large models such as LLaMA3.1-70B can be quantized on a single A6000 GPU. In contrast, efficient finetuning–based methods (Shao et al., 2023; Dettmers et al., 2023b, a; Chee et al., 2023; Li et al., 2023; Xu et al., 2023; Chen et al., 2024b) require partial retraining, which increases time cost but often mitigates VRAM pressure by freezing most parameters, training only a small subset, or performing quantization in batches. This strategy allows quantization to be completed within limited video memory. However, while prior studies have primarily focused on reducing GPU memory consumption, the CPU memory overhead has been largely neglected. Existing quantization frameworks typically require loading the entire model into CPU memory during processing, which poses a critical bottleneck for large-scale models. For instance, quantizing a LLaMA 3.1–70B model requires at least 140 GB of CPU RAM, and the actual usage is often much higher, exceeding the capacity of most standard servers. In this paper, we integrate our proposed Hierarchical Linear Quantization with efficient quantization techniques to enhance model accuracy while maintaining low GPU memory usage, reduced CPU memory footprint, and short quantization time. For example, our framework can quantize the Llama 3.1-70B model using 2 RTX 4090 GPU (48GB VRAM) and only 64GB of CPU RAM, significantly reducing both GPU and CPU memory requirements.
Although weight quantization reduces memory footprint, traditional methods require dequantization to higher precision (8-bit/FP16) for computation, introducing significant overhead that can paradoxically slow down inference at low bit-widths. Recent advances (Wei et al., 2025; Park et al., 2025b, 2022) address this by replacing standard GEMM with lookup table (LUT)-based operations that implement generalized FP-INT multiplication, eliminating dequantization while achieving both linear latency reduction with bit-width and improved energy efficiency. FLGLUT (Park et al., 2025b) accelerates these operations on GPUs and T-MAC (Wei et al., 2025) optimizes for CPUs via SIMD instructions. We enhance this paradigm through a pure C++ kernel redesign that specifically supports our novel Hierarchical Linear Quantization format, maintaining cross-platform compatibility while optimizing for our method’s unique requirements.
@@ 2.2 LUT-Based GEMM
Figure 2(c) illustrates another GEMM paradigm termed Bit-serial LUT-based GEMM. Unlike the scheme in Figure 2(b), this method utilizes the LUT to store all possible dot products between an activation vector and single-bit weights. The detail computational procedure is shown in Figure 3. First, a $q$-bit quantized weight matrix $\mathrm{W}_{int}$ is decomposed into $q$ single-bit matrices $\{W_{0},W_{1},...,W_{q-1}\}$ offline, where each element is either 0 or 1, representing the respective bit planes of the original weights. For example, for integer values (9, 7, 6, 3) with binary representations (1001, 0111, 0110, 0011), the matrix for the lowest bit is (1, 1, 0, 1), and the matrix for the highest bit is (1, 0, 0, 0). This decomposition is performed offline, incurring no runtime overhead. During inference, for an activation vector of the group size $g$, the system precomputes the dot products between this activation vector and all $2^{g}$ possible combinations of single-bit weights, storing the results in the LUT. Thus, the original matrix computation requiring high-precision multiply-accumulate operations is simplified into highly efficient table lookups followed by summation. This paradigm has been demonstrated to offer high computational efficiency and energy efficiency (Park et al., 2025b; Wei et al., 2025). For example, FIGLUT (Park et al., 2025b) optimized the table structure for GPU architectures to avoid bank conflicts, while T-MAC leveraged CPU vectorized lookup instructions (AVX2/NEON) to enable efficient LUT operations on CPUs.
@@ 3.1 Bell-Shaped Distribution of Weights
As noted in prior studies (Kim et al., 2023; Dettmers et al., 2023a; Zhao & Yuan, 2025), weights often exhibit a bell-shaped distribution. Uniform quantization, which maps weights to a uniformly spaced grid, struggles to accurately capture such distributions, especially under low-bit settings. Figure 4 shows that the weight distribution of an out_channel in the up_projection layer of the LLaMA3.1-8B model closely follows a Gaussian pattern, with values heavily concentrated near zero. This characteristic, commonly observed across various models, highlights the inherent limitation of uniform quantization in effectively representing natural weight distributions. Consequently, non-uniform quantization strategies are often necessary to achieve higher fidelity in low-bit compression.
Figure 4: (Top) The weight distribution of one output channel in a up_proj of LLaMA3.1-8B. (Bottom) MSE loss for weight comparsion between Uniform and HLQ.
@@ 3.2 Optimization for Bit-Serial LUT-Based GEMM
For ordinary LUT-based GEMM algorithms shown in figure 2(b), previous efforts such as SqueezeLLM (Kim et al., 2023) and GANQ (Zhao & Yuan, 2025) have achieved significant improvements in algorithmic performance for 3-bit and 4-bit quantization, along with the development of highly efficient CUDA kernels for accelerated inference. However, these implementations still lack support for CPU devices. Regarding Bit-serial LUT-Based GEMM, existing works including T-MAC (Wei et al., 2025), LUT-GEMM (Park et al., 2022), and FIGULT (Park et al., 2025b) have primarily focused on optimizing kernel design to support efficient uniform quantization schemes, yet they do not address algorithmic enhancements for improving quantization accuracy. Contemporary work like AnyBCQ (Park et al., 2025a) has made certain explorations, focusing on designing multi-precision models, with its 2-bit models achieving limited performance and primarily targeting model inference on GPUs. This limitations restricts the broader application of such kernels in cpu-based edge devices. To bridge this gap, our work proposes using HLQ to boost quantization accuracy at the algorithm level, while leveraging Bit-Serial LUT-Based GEMM to enable efficient inference on CPU devices.